OpenStudy (anonymous):

If vector A⃗ has components Ax and Ay and makes an angle θ with the +x axis, then A) A=Ax+Ay (where Ais the magnitude of A⃗ ) B) θ =AyAx C) cosθ = AxA2x+A2y√ D) tanθ =AxAy

OpenStudy (anonymous):

then what : ) ?

OpenStudy (anonymous):

A) A=Ax+Ay (where Ais the magnitude of A⃗ ) B) θ =AyAx C) cosθ = AxA2x+A2y√ D) tanθ =AxAy

OpenStudy (anonymous):

Have you typed the options accurately ? I think C must have a typo error, and maybe D as well

OpenStudy (anonymous):

C) cosθ = \[\frac{ Ax }{ \sqrt{A \frac{ 2 }{ x } + A \frac{ 2 }{ y }} }\]

OpenStudy (anonymous):

D) tanθ = \[\frac{ Ax }{ Ay }\]

OpenStudy (anonymous):

There are a few things you need to know to answer this question, nothing too complicated though.

OpenStudy (anonymous):

like, how are the trig functions cos and tan defined, and how is the length of a vector calculated from its components and it would help if you draw yourself a diagram of the vector and its components

OpenStudy (anonymous):

Im pretty sure tanθ = is y/x so d would be wrong